A ball is spinning in a vertical circle at the end of a string that is 3.0 m long . If the ball has a mass of 2.5 kg and moves at constant speed of 6.0m/s determine the tension in the string when the ball is at top of the loop. Determine the tension in the string when the ball is the bottom of the loop.


1
Expert's answer
2021-12-15T16:22:34-0500

The tension in the string when the ball is at top of the loop


"mg+T=ma\to mg+T=m\cdot\frac{v^2}{r}\to T=m\cdot\frac{v^2}{r}-mg="


"=2.5\cdot\frac{6^2}{3}-2.5\cdot9.8=5.5\ (N)"


The tension in the string when the ball is the bottom of the loop


"mg-T=-ma\to mg-T=-m\cdot\frac{v^2}{r}\to T=m\cdot\frac{v^2}{r}+mg="


"=2.5\cdot\frac{6^2}{3}+2.5\cdot9.8=54.5\ (N)"